OpenStudy (anonymous):

How many years older will you be 1.20 billion seconds from now? assume you are 20 yrs old

OpenStudy (anonymous):

1.2 billion years equal 38,05 years. that makes 58,05 years old

OpenStudy (anonymous):

1.2 billion seconds equals 38.0265 years. That would make you 58 years old!

OpenStudy (anonymous):

1 year = 365 days 1 day = 24 hours 1 hour = 60 minutes 1 minute = 60 seconds You multiply by a chain of conversion factors, to convert from 1.20 billion seconds to years. These conversion factors work because they are equal to 1 - their numerators and denominators cancel out. \[1.20\times10^9\sec=(1.20\times10^9 \sec)(\frac{1 \min}{60 \sec})(\frac{1\ hr}{60 \min})(\frac{1\ day}{24\ hr})(\frac{1\ yr}{365\ day})=38.05\ yr\\38.05\ yr+20\ yr=58.05\ yr\]